4. a) John LTD is a listed company on the Ghana Stock Exchange Market and showed the following performance. The following was made available to you.
Current market price per share (as at 31/12/2017)=$10
Dividend per share 2017=$1
Expected growth rate returns=27%
The government treasury bill rate=24%
The beta factor of john ltd=1.4
Required
i)what is the estimated cost of equity using the dividend growth model?
ii)what is the estimated cost of equity using the Capital Assets Pricing model?
1
Expert's answer
2020-11-17T06:38:57-0500

i) find the formula:

"Ke=\frac{D}{P}+g=\frac{1}{10}+0.27=0.37"

ii) find the formula:

"ri=rf+\beta (rm-rf=24+1.4(27-24)=28.2"